OpenStudy (kewlgeek555):

I have Eight Algebra Questions about Solving Systems of Equations Algebraically That I Would Like You To Check if They Are Correct. Please?

OpenStudy (kewlgeek555):

I. Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -2 a. x = 0 b. x = 2 c. There is no x value as there is no solution. d. x can be any value as there is an infinite number of solutions.

OpenStudy (kewlgeek555):

I chose C, There is no x-value as there is no solution.

OpenStudy (anonymous):

D. Only because Y is not defined. Y can be any number, resulting in x being any number.

OpenStudy (anonymous):

If that makes sense. It been awhile since I was in that math. Calc has changed me.

OpenStudy (kewlgeek555):

Thank you @mostwanting Can you please see if you can see if I got the right answers on the other seven questions.

OpenStudy (anonymous):

Sure

OpenStudy (kewlgeek555):

Thank you.

OpenStudy (kewlgeek555):

Question 2 (Multiple Choice Worth 1 points) (05.02) If you were solving a system of equations and you came to a statement like 1 = 3, what do you know about the solution to the system? The solution is (1, 3) The solution is x = 1 and y = 3 There is no solution There are infinitely many solutions

OpenStudy (kewlgeek555):

I chose C.

OpenStudy (anonymous):

Agreed. Next.

OpenStudy (kewlgeek555):

Okay. Sorry for the wait, hold on.

OpenStudy (anonymous):

I'll be here, lol

OpenStudy (kewlgeek555):

The art museum had a total of 224 visitors on Tuesday. Visitors older than 18 paid $12 for admission. Visitors 18 years or younger paid $8 for admission. The museum collected a total of $2,520. How many visitors were 18 or younger? 42 56 182 194 I dont know this one. D;

OpenStudy (anonymous):

Ok, let me write it out on paper, cant do this in my head, but I do know how to do it. Just give me a min or two.

OpenStudy (anonymous):

If I give you an equation, you think You can figure it out?  With this, and I believe you know how to do it with this info, you should solve with x. I wont give you this answer, only because I think you didnt know how to set it up.

OpenStudy (kewlgeek555):

Okay, thanks! That is what I had a trouble with - setting it up! I didnt do the 8x + 12y thingie that you did, let me solve it.

OpenStudy (anonymous):

Ok, good.

OpenStudy (kewlgeek555):

Okay, um...I got 315 solving for x...

OpenStudy (kewlgeek555):

This is what I did:\[8x+12y=2520\]\[8x+12(0)=2520\]\[8x+0=2520\]\[8x=2520\]\[315.\]

OpenStudy (anonymous):

I solved for Y, than plugged it in to the 2nd equation and got 14... I didnt check my own answer. Let me think again.

OpenStudy (anonymous):

Post the next question while I figure what I did wrong

OpenStudy (kewlgeek555):

Oh, okay. Question 5 (Multiple Choice Worth 1 points) (05.02) Solve the system of equations. 2x - 5y = 3 x - 3y = 1 (4, 1) (5, 2) (7, 2) (9, 3) I chose A

OpenStudy (anonymous):

Agreed

OpenStudy (kewlgeek555):

I got -91 as my answer for the other one.

OpenStudy (anonymous):

Hmmm. Do the next. I either get negatives or 315 and stuff. I'll try to rewrite my equation

OpenStudy (kewlgeek555):

Okay. Solve the following system of equations: -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3) I chose D.

OpenStudy (anonymous):

That good.

OpenStudy (kewlgeek555):

Okay. Solve the following system of equations: x - 2y = 14 x + 3y = 9 (1, 12) (-1, -12) (12, -1) (12, 1)

OpenStudy (anonymous):

C. I'll give you that one

OpenStudy (kewlgeek555):

Okay, thanks. Oh, I forgot to tell you what I chose. I chose C actually. Lol. I have two more that I HAVENT answered Theyre the same subject.

OpenStudy (kewlgeek555):

Ill be back like in 20 minutes.

OpenStudy (anonymous):

Ok. I'll help some others and look up the other problem thats a pain.

OpenStudy (kewlgeek555):

Okay, I'm back.

OpenStudy (kewlgeek555):

Here is one of the two questions I cannot seem to figure out. Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18 a. 12x + 4y = -12 15x = -30 b. 8x + 4y = -4 14x = -10 c. 6x - 8y = -36 -6y = -42 d. 6x - y = -15 3y = 9

OpenStudy (anonymous):

did u ever gdt the answer?

OpenStudy (anonymous):

ya @kewlgeek555

OpenStudy (anonymous):

what was the answer then?

OpenStudy (anonymous):

If x can be any number, how many solutions are there for the equation? y = 4x – 1 A. There are two solutions. B. There is no solution. C. There are many solutions. D. There is only one solution.